OpenStudy (anonymous):

how can i sketch the curve z=k for the specified values of k? if z = x^2 + y ; k = -2, -1, 0, 1, 2

OpenStudy (anonymous):

yes but actually our topic in that is about contour plotting..

OpenStudy (stamp):

how can i sketch the curve z=k for the specified values of k? if z = x^2 + y ; k = -2, -1, 0, 1, 2 -2 = x^2 + y -1 = x^2 + y 0 = x^2 + y 1 = x^2 + y 2 = x^2 + y sketch = http://www.wolframalpha.com/input/?i=plot+-2+%3D+x^2+%2B+y+and+-1+%3D+x^2+%2B+y+and+0+%3D+x^2+%2B+y+and+1+%3D+x^2+%2B+y+and+2+%3D+x^2+%2B+y+from+-3+to+3

OpenStudy (anonymous):

how can i plot that equations?

OpenStudy (anonymous):

how can i plot those equations? -2, -1, 0, 1, 2 -2 = x^2 + y -1 = x^2 + y 0 = x^2 + y 1 = x^2 + y 2 = x^2 + y

OpenStudy (stamp):

-2 = x^2 + y y = -2 - x^2 It is a parabola

OpenStudy (anonymous):

ah..ok tnx but how can i know that it is a parabola?

OpenStudy (stamp):

@jedai17 Because hopefully you took an Algebra II class before taking Calculus III and you remember studying conic sections and how to graph parabolas if not, here is a review http://tutorial.math.lamar.edu/Classes/Alg/Parabolas.aspx

OpenStudy (anonymous):

im sorry i forgot about it..

OpenStudy (stamp):

that is why i provided a link for you to review

OpenStudy (anonymous):

ok thankyou i will review it..

OpenStudy (anonymous):

btw it it the same if z = k in z = x^2 +9y^2 ; k = 0, 1, 2, 3, 4 ??

OpenStudy (stamp):

Yes. Except k = x^2 + 9y^2 will be an ellipse rather than a parabola.

OpenStudy (anonymous):

ah ok...thank you so much :)

OpenStudy (anonymous):

how to find the vertex of this equation? y = -2 - x^2 ?

OpenStudy (stamp):

\[y=-x^2-2\] vertex x value = -b/2a \[ax^2+bx+c\] a = -2, b = 0, c = -2 -b/2a = - ( 0 ) / 2( - 2 ) = 0 So vertex occurs at x = 0, or ordered pair (0, -2)

OpenStudy (stamp):

I made it from -3 to 3. It looks ugly if it is from 0.000001 to - 3250123989. It is an arbitrary domain I chose to give you a full picture of the graph.

OpenStudy (anonymous):

ah ok..tnx

OpenStudy (anonymous):

in this y=−x2−2 vertex x value = -b/2a ax2+bx+c a = -2, b = 0, c = -2 -b/2a = - ( 0 ) / 2( - 2 ) = 0 So vertex occurs at x = 0, or ordered pair (0, -2) why the value of a is -2? i think it is -1..

OpenStudy (stamp):

It is a = -1. My mistake.

OpenStudy (stamp):

But since it is -b / 2a, and b = 0, it is fortunately irrelevant in this case.

OpenStudy (anonymous):

its ok thanks again

OpenStudy (anonymous):

what do u mean by irrelevant?

OpenStudy (stamp):

If -b / 2a, and b = 0, does it matter what a is ? No. 0 / anything = 0. So what is the difference between 0 / 2(-2) and 0 / 2(-1) They both equal 0. Even though I made a mistake on my a value, we were fortunately ok because a did not effect the outcome. irrelevant - Not connected with or relevant to something.

OpenStudy (anonymous):

ah ok thank you so much for this infos.. :)